(7x^2/7)-3=25. solve for x
Answer by TimothyLamb(4379)

---
(7x^2/7) - 3 = 25
(7/7)x^2 + (0)x - 28 = 0
x^2 + (0)x - 28 = 0
---
the above quadratic equation is in standard form, with a=1, b=0, and c=-28
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 0 -28
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the two real roots (i.e. the two x-intercepts), of the quadratic are:
x = 5.29150262
x = -5.29150262
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php